Income Or Output Y Consumption Expenditure         C Investment Expenditure       I Government Expenditure          G...

Income

Or

Output

Y

Consumption

Expenditure

        C

Investment

Expenditure

      I

Government

Expenditure

         G

Net export

Expenditure

       NX

$4,000

3,925

100

100

25

4,100

4,000

100

100

25

4,200

4,075

100

100

25

4,300

4,150

100

100

25

4,400

4,225

100

100

25

4,500

4,300

100

100

25

4,600

4,375

100

100

25

4,700

4,450

100

100

25

4,800

4,525

100

100

25

4,900

4,600

100

100

25

5,000

4,675

100

100

25

a) Determine equilibrium level of income?

b) Calculate expenditure multiplier?

c) If government expenditure grows to $150, what will be the new equilibrium level of income?

d) If government expenditure falls to $50, what will be the new equilibrium level of income?

e) Suppose a tax cut of $50 is given with the information given in question1, what will be the new equilibrium level of income after the tax cut?

2. Calculate MPC, MPS and the Multiplier if consumption expenditure increases by $4,250 as a result of increase in income from $40,000 to $45,000.

3. Calculate GDP loss if equilibrium level of GDP is $10,000, unemployment rate 9.8%, and the MPC is 0.90.

Show it with help of a graph taking AE on the vertical axis and income on the horizontal axis.
Answer #1

Income Consumption Investment Government Net exportAggregate Expenditure Expenditure Expenditure Expenditure Expenditure Output, Y $4,000 4.100 4,200 4,300 4,400 4,500 4,600 4,700 4,800 4,900 5,000 3,925 4,000 4,075 4,150 4,225 4,300 4,375 4,450 4,525 4,600 4,675 100 100 100 100 100 100 100 100 100 100 100 100 100 100 100 100 100 100 100 100 100 100 25 25 25 25 25 25 25 25 25 25 25 4,150 4,225 4,300 4,375 4,450 4,525 4,600 4,675 4,750 4,825 4,900

a) Equilibrium level of income = 4600 because Y = AD or Y = AE

b) Expenditure multiplier = 1/ ( 1 - MPC)

MPC = Change in Consumption / Change in Output = 75/100 = 0.75

Change in consumption is 75 at each output.

Change in output is fixed at 100 units.

Expenditure multiplier = 1/(1 - 0.75) = 1/0.25 = 4

c) New equilibrium level of income is 4800.

Income Consumption Investment Government Net exportAggregate Expenditure Expenditure Expenditure Expenditure Expenditure Output, Y $4,000 4.100 4,200 4,300 4,400 4,500 4,600 4,700 4,800 4,900 5,000 3,925 4,000 4,075 4,150 4,225 4,300 4,375 4,450 4,525 4,600 4,675 100 100 100 100 100 100 100 100 100 100 100 150 150 150 150 150 150 150 150 150 150 150 25 25 25 25 25 25 25 25 25 25 25 4,200 4,275 4,350 4.425 4,500 4,575 4,650 4,725 4,800 4,875 4,950

d) New equilibrium level of income = 4400

Income Consumption Investment Government Net exportAggregate Expenditure Expenditure Expenditure Expenditure Expenditure Output, Y $4,000 4.100 4,200 4,300 4,400 4,500 4,600 4,700 4,800 4,900 5,000 3,925 4,000 4,075 4,150 4,225 4,300 4,375 4,450 4,525 4,600 4,675 100 100 100 100 100 100 100 100 100 100 100 50 50 50 50 50 50 50 50 50 50 50 25 25 25 25 25 25 25 25 25 25 25 4.100 4,175 4,250 4,325 4,400 4,475 4,550 4,625 4,700 4,775 4,850

2. MPC = Change in C / Change in Y = 4250/(45000 - 40000) = 4250/5000 = 0.85

MPS = 1 - MPC = 1 - 0.85 = 0.15

Multiplier = 1/MPS = 1/0.15 = 6.67